# Improving Probabilistic Diffusion Models With Optimal Covariance Matching

### v2

@@ 5.3 Image Modelling with Latent Diffusion Models
6 Related Work and Future Directions
In addition to the covariance estimation methods discussed in Section 4, there are other approaches to accelerate sampling in diffusion models. One approach involves using faster numerical solvers for differential equations with continuous timesteps (Jolicoeur-Martineau et al., 2021; Liu et al., 2022; Lu et al., 2022). Another strategy, inspired by Schrödinger bridge (Wang et al., 2021; De Bortoli et al., 2021), is to introduce a nonlinear, trainable forward diffusion process. Additionally, combining diffusion models with GANs enhances GAN sample quality while speeding up diffusion model sampling (Xiao et al., 2021; Wang et al., 2022).
Recently, distillation techniques have gained popularity, achieving state-of-the-art in one-step generation (Zhou et al., 2024). There are two prominent types of distillation methods. The first is trajectory distillation (Salimans & Ho, 2022; Berthelot et al., 2023; Song et al., 2023; Heek et al., 2024; Kim et al., 2023; Li & He, 2024), which focuses on accelerating the process of solving differential equations. The second type involves distillation techniques that utilize a one-step implicit latent variable model as the student model (Luo et al., 2024; Salimans et al., 2024; Xie et al., 2024; Zhou et al., 2024) and distills the diffusion process into the student model by minimizing the spread divergence family (Zhang et al., 2020; 2019) through score estimation (Poole et al., 2022; Wang et al., 2024).
Beyond image modelling, our method can be straightforwardly applied to accelerate large-scale video diffusion models (Blattmann et al., 2023; Chen et al., 2023), which are based on latent diffusion models. Furthermore, a recent study (Zhao et al., 2024) shows that covariance estimation is crucial in addressing the image-leakage issue in image-to-video generation problems, which presents another important application of our method that we will explore in future work.
@@ 7 Conclusion
In this paper, we proposed a novel method for learning the optimal diagonal covariance of the denoising distribution, which offers improved accuracy compared to other covariance learning approaches. Our results demonstrate that enhanced covariance estimation leads to better generation quality and diversity, all while reducing the number of required generation steps. We validated the scalability of our method across different problem domains and scales. Additionally, we discussed its connection to various acceleration techniques and highlighted several promising application areas for future exploration.
Acknowldgement
We want to thank Wenlin Chen for the useful discussions. MZ and DB acknowledge funding from the Cisco Centre of Excellence. ZO is supported by the Lee Family Scholarship.
@@ A.3 Connection to SN-DDPM
where $\epsilon=(\tilde{x}-\sqrt{\alpha}x)/\sqrt{\beta}$. Now we can apply Equations 32 and 33 to establish the connection between OCM-DDPM and SN-DDPM, as demonstrated in the following theorem.
@@ B.1 Details of Model Architectures
ImageNet 256x156 $\mathrm{AdaLN}$+$\mathrm{Linear}$ $\mathrm{AdaLN}$+$\mathrm{Linear}$
@@ Appendix C Additional Experimental Results
In this section, we present additional experimental results for comparison. First, we present an additional likelihood comparison between I-DDPM, NPR-DDPM, and OCM-DDPM in Section C.1 and discuss the performance variance in Section C.2. Next, we provide further experimental results on latent diffusion models in Section C.3 and conduct qualitative studies by showcasing samples generated by our models in Section C.4.
C.1 Additional Likelihood Comparison
Table 10: Upper bound on the negative log-likelihood (bits/dim) on the ImageNet 256x256 dataset.
@@ C.1 Additional Likelihood Comparison
C.2 Mean and Variance of Performance
@@
C.3 More Results on Latent Diffusion Models
@@ C.4 Generated Samples
  • In Figures 7, 8, 9 and 10, we visualize the generated samples of our models using different number of sampling steps on CIFAR10 (LS), CIFAR10 (CS), CelebA 64x64, and ImageNet 64x64, respectively (refer to Table 5).
  • In Figure 11, we visualize the generated samples on ImageNet 256x256, using 10 sampling steps with different CFG coefficients. (refer to Figures 4 and 11).

### v4

@@ 7 Conclusion
In this paper, we proposed a new method for learning the diagonal covariance of the denoising distribution, which offers improved accuracy compared to other covariance learning approaches. Our results demonstrate that enhanced covariance estimation leads to better generation quality and diversity, all while reducing the number of required generation steps. We validated the scalability of our method across different problem domains and scales and discussed its connection to various acceleration techniques and highlighted several promising application areas for future exploration.
8 Acknowledgments
ZO is supported by the Lee Family Scholarship. MZ and DB acknowledge funding from AI Hub in Generative Models, under grant EP/Y028805/1 and funding from the Cisco Centre of Excellence. We want to thank Wenlin Chen for the useful discussions.
@@ A.3 Connection to SN-DDPM
where $\epsilon=(\tilde{x}-\sqrt{\alpha}x)/\sqrt{\beta}$. Now we can apply Equations 31 and 32 to establish the connection between OCM-DDPM and SN-DDPM, as demonstrated in the following theorem.
@@ B.1 Details of Model Architectures
ImageNet 256x256 $\mathrm{AdaLN}$+$\mathrm{Linear}$ $\mathrm{AdaLN}$+$\mathrm{Linear}$
@@ B.2 Details of Training, Inference, and Evaluation
(a) Cov Error with True Score
(b) Cov Error with Learned Score
(c) DDPM MMD v.s. Steps
(d) DDIM MMD v.s. Steps
Figure 4: Comparisons of different covariance estimation methods based on estimation error and sample generation quality. Figures (a) and (b) show the mean square error of the estimated diagonal covariance under the assumptions: (a) access to the true score, and (b) learned score of the data distribution at various noise levels. Figures (c) and (d) present the MMD evaluation against the total sampling steps in the DDPM (c) and DDIM (d) settings. We can find the proposed OCM method can achieve the lowest estimation error and consistently outperform other baseline methods when fewer generation steps are applied.
@@ Appendix C Additional Experimental Results
In this section, we present additional experimental results for comparison. First, we present an additional toy experiment in Section C.1 and likelihood comparison between I-DDPM, NPR-DDPM, and OCM-DDPM in Section C.2 and present results using different numbers of Rademacher samples in Section C.3. We also analyze the performance variance in Section C.4. Additionally, we include further experimental results on latent diffusion models in Section C.5 and conduct qualitative studies by showcasing samples generated by our models in Section C.6.
C.1 Additional Toy Demonstration
To further verify the effectiveness of our method, we include an additional toy example. In this case, we consider another two-dimensional mixture of nine Gaussians (MoG) with means located at $\{-3,0,3\}\otimes\{-3,0,3\}$ and a standard deviation of $\sigma=0.1$. To assess different approaches, we first learn the covariance using the true scores. Specifically, we train the covariance networks for these methods over 50,000 iterations with a learning rate of $0.001$ using an Adam optimizer (Kingma & Ba, 2014). Figure 4(a) shows the $L_{2}$ error of the estimated diagonal covariance $\Sigma_{t{-}1}(x_{t})$, and our method, OCM-DDPM, consistently achieves the lowest error compared to the other methods. In practice, the true score is not accessible. Therefore, we also conduct a comparison using a score function learned with DSM. We then apply different covariance estimation methods under the same settings. In Figure 4(b), we plot the same $L_{2}$ error for the learned score setting and find that our method achieves the lowest error for all $t$ values.
We further use the learned score and the covariance estimated by different methods to conduct DDPM and DDIM sampling for this MoG problem. Figures 4(c) and 4(d) presents the MMD comparison across various methods. It shows that our method outperforms the baselines when the total time step is small, demonstrating the importance of accurate covariance estimation in the context of diffusion acceleration. We also include two methods utilizing the true diagonal and full covariance as benchmarks, representing the best achievable performance. Notably, these two methods exhibit similar performance because the MoG used in this setting has symmetric components, in which the covariance is dominated by the diagonal entries.
C.2 Additional Likelihood Comparison
Table 10: Upper bound on the negative log-likelihood (bits/dim) on the ImageNet 64x64 dataset.
@@ C.2 Additional Likelihood Comparison
C.3 Impact of the Numbers of Rademacher Samples
Table 11: Results of OCM-DDPM on CIFAR10 (CS) with varying numbers of Rademacher Samples.
FID $\downarrow$ NLL (%) $\uparrow$
$K=$ 10 25 50 100 10 25 50 100
$M=1$ 14.32 5.54 4.10 3.84 4.99 4.34 3.99 3.76
$M=3$ 14.18 5.51 4.11 3.82 4.99 4.34 3.99 3.76
$M=5$ 14.17 5.51 4.11 3.82 4.99 4.34 3.99 3.76
$M=10$ 14.16 5.51 4.11 3.82 4.99 4.34 3.99 3.76
In this section, we include an ablation study examining the effect of varying the number of Rademacher samples $M$. Specifically, we conduct experiments on CIFAR10 (CS) using the proposed OCM-DDPM method and evaluate performance based on FID and NLL. The empirical results, presented in Table 11, indicate that a larger $M$ (e.g., $M=3$) can give a small improvement in FID. This improvement is likely due to the reduced gradient estimation variance during training with a larger $M$. However, in most cases, different values of $M$ yield consistent performance. This is practically desirable as, in practice, setting $M=1$ allows for efficient training while maintaining strong performance
C.4 Mean and Variance of Performance
@@
C.5 More Results on Latent Diffusion Models
@@ C.6 Generated Samples
  • In Figures 8, 9, 10 and 11, we visualize the generated samples of our models using different number of sampling steps on CIFAR10 (LS), CIFAR10 (CS), CelebA 64x64, and ImageNet 64x64, respectively (refer to Table 5).
  • In Figure 12, we visualize the generated samples on ImageNet 256x256, using 10 sampling steps with different CFG coefficients. (refer to Figures 3 and 12).